OpenStudy (chrisbax):

what is a slope of a line with a rise and a run of 4?

candycove (candycove):

http://openstudy.com/study#/groups/Mathematics

rebeccaxhawaii (rebeccaxhawaii):

\(\style{background:black;padding:5px;border:blue 3px ridge;text-shadow:blue 1px 1px 0px;font-size:34px;}{\huge\mathbb{\color{darkcyan}{WELCOME\color{cyan}{~TO}~OPENSTUDY}}}\) Hey i have made this to hopefully answer any questions you may have. But if it doesn't feel free to pm me c: http://openstudy.com/users/rebeccaxhawaii#/updates/56f5205ce4b07a8c82287f75 Good Luck with your studies! @chrisbax \(\Huge\heartsuit\) \[\frac{ rise }{ run }=\frac{ 4 }{ 4 }\]

rebeccaxhawaii (rebeccaxhawaii):

Rise over run is the slope - therefore simplify 4/4 to get your slope.

ILovePuppiesLol (ilovepuppieslol):

Welcome to Openstudy :) @chrisbax